# When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs

### v2

@@ 1 Introduction
Self-correction is a popular approach to improve responses from large language models (LLMs) by refining them using LLMs during inference (Bai et al., 2022; Madaan et al., 2023). Extensive studies on self-correction have been conducted in various tasks, including arithmetic reasoning, code generation, and question answering (Gao et al., 2023; Shinn et al., 2023). The simplest approach of self-correction prompts LLMs to provide feedback on their own responses and refine the responses using the feedback (Huang et al., 2024a), under the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022). As in Figure 1, self-correction has also been studied using additional information for improving feedback, including external tools such as code interpreters (Chen et al., 2024d; Gou et al., 2024), external knowledge retrieved via web search (Gao et al., 2023; Jiang et al., 2023b), or fine-tuning (Welleck et al., 2023; Ye et al., 2023). However, recent studies also report negative results indicating that LLMs cannot self-correct (Huang et al., 2024a; Gou et al., 2024; Li et al., 2024b) or even self-detect (Chen and Shu, 2024; Tyen et al., 2024; Hong et al., 2024; Jiang et al., 2024; Kamoi et al., 2024) their own mistakes at least in certain conditions. These conflicting observations indicate that further analysis of self-correction is needed.
@@
Self-Debug (2024d) Fair-Asym. GPT-3.5-T, GPT-4 Code Interpreter Text-to-Code
@@
Tools , color=Red!100, fill=Purple!15, thick, text=black, l=5pt [ Fair , color=Black!100, fill=White!15, thick, text=black [ Code Interpreter or Compiler , color=Red!100, fill=Purple!15, thick, text=black [ Self-Debug (2024d), CRITIC (2024),
@@
Tools , color=Red!100, fill=Orange!15, thick, text=black, l=5pt [ Fair , color=Black!100, fill=White!15, thick, text=black [ Code Interpreter or Compiler , color=Red!100,, fill=Orange!15, thick, text=black, [ Self-Edit (2023b, SFT, PyCodeGPT) , leaf, color=Red!50, fill=Orange!15, thick, text=black ] ] [ Proof Assistant , color=Red!100, fill=Orange!15, thick, text=black, [ Baldur (2023, SFT) , leaf, color=Red!50, fill=Orange!15, thick, text=black ] ] ] ] ] ] [ Cross-Model
@@
Tools , color=Red!100, fill=Green!15, thick, text=black, l=5pt [ Code Interpreter or Compiler, color=Red!100, fill=Green!15, thick, text=black, text width=10em[ CodeRL (2022, RL), Self-Edit (2023b, SFT, GPT-3) , leaf, color=Red!50, fill=Green!15, thick, text=black, text width=20em] ] ] ] [ Multi-Agent
@@ 2.1 Frameworks
Cross-model correction generates feedback or refines the responses using models different from the model that generates initial responses. Cross-model correction has been mostly studied in the settings of correcting mistakes of large proprietary LLMs using small fine-tuned models (Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024) or multi-agent debate of multiple models with similar capabilities (Liang et al., 2023; Li et al., 2023; Cohen et al., 2023; Du et al., 2023; Zhang et al., 2023a; Chen et al., 2024b; Chan et al., 2024; Wang et al., 2024a).
@@ 2.2 Sources of Feedback
Intrinsic self-correction prompts LLMs to generate feedback on their own responses. Prompting strategies include simple zero-shot or few-shot prompts (Madaan et al., 2023; Kim et al., 2023), decomposing the responses (Dhuliawala et al., 2023), and evaluating confidence (Varshney et al., 2023; Jiang et al., 2023b; Wu et al., 2024).
@@
Self-correction often relies on external information, including external tools such as code executors (Jiang et al., 2023a; Gou et al., 2024; Chen et al., 2024d; Stengel-Eskin et al., 2024), symbolic reasoners (Pan et al., 2023), proof assistant (First et al., 2023), or task-specific metrics (Xu et al., 2023), external knowledge from search engines (Jiang et al., 2023b; Gao et al., 2023; Zhao et al., 2023), Wikipedia (Yu et al., 2023; Zhao et al., 2023), or other corpora (Peng et al., 2023; Zhao et al., 2023), oracle information such as ground-truth answers (Kim et al., 2023; Shinn et al., 2023), human feedback (Chen et al., 2024a), or stronger models (Zhang et al., 2024).
@@
Models fine-tuned for self-correction are another source of feedback, which are trained via supervised fine-tuning (Welleck et al., 2023; Ye et al., 2023; First et al., 2023; Paul et al., 2024; Han et al., 2024) or reinforcement learning (Le et al., 2022; Akyurek et al., 2023).
@@ 2.3 Tasks
Self-correction has been studied in various tasks, including Reasoning: arithmetic reasoning (Madaan et al., 2023; Nathani et al., 2023; Gou et al., 2024), code generation (Jiang et al., 2023a; Charalambous et al., 2023; Gou et al., 2024; Chen et al., 2024d; Olausson et al., 2024), proof generation (First et al., 2023), logical reasoning (Pan et al., 2023), Knowledge: closed-book QA (Shinn et al., 2023; Gao et al., 2023; Jiang et al., 2023b; Gou et al., 2024), Context-based Generation: dialogue generation (Madaan et al., 2023; Peng et al., 2023), text summarization (Saunders et al., 2022), Open-ended Generation: conditional text generation (Ye et al., 2023; Schick et al., 2023), story generation (Yang et al., 2022b), detoxification (Schick et al., 2021; Bai et al., 2022; Gou et al., 2024; Phute et al., 2024), Others: machine translation (Chen et al., 2023b; Raunak et al., 2023; Ki and Carpuat, 2024), information retrieval (Gero et al., 2023), vision language tasks (Yin et al., 2023; Ge et al., 2023; Zhou et al., 2024; Lee et al., 2024; Huang et al., 2024b; Liu et al., 2024), and prompt optimization (Pryzant et al., 2023; Mehrabi et al., 2023; Yang et al., 2024).
@@ 3.2 Frameworks for Verifying RQs
Realistic frameworks can be categorized by whether they use the best-possible initial responses. Fair self-correction represents frameworks that refine the best-possible initial responses. (1) Intrinsic self-correction (Huang et al., 2024a) uses the same model and information for initial response generation and self-correction. Intrinsic self-correction can be used to assess [RQ1] whether LLMs can self-correct based solely on their inherent capabilities. (2) Fair-asymmetric self-correction uses additional information for self-correction, but also uses information to improve initial response generation as much as possible. For example, self-correction with code interpreters (Chen et al., 2024d; Gou et al., 2024) is not intrinsic but fair because we cannot easily use code interpreters to directly improve the initial response generation. Fair-asymmetric self-correction can be used to evaluate [RQ2] whether LLMs can self-correct the best-possible initial responses using external information. Unfair self-correction (or unfair-asymmetric self-correction) represents frameworks that are practical but do not use the best-possible initial responses. For example, methods that use search engines only for self-correction (Gao et al., 2023; Yu et al., 2023) are unfair because they can use search engines to directly improve the initial response generation. Unfair self-correction can evaluate [RQ3] whether the final responses from self-correction outperform other methods but cannot evaluate [RQ2] whether self-correction can improve the best-possible initial responses.
@@ 4 Self-Correction with Prompting
However, recent studies report that intrinsic self-correction does not improve or even degrade the performance in tasks such as arithmetic reasoning, closed-book QA (Huang et al., 2024a; Gou et al., 2024), code generation (Gou et al., 2024; Olausson et al., 2024), plan generation (Valmeekam et al., 2023), and graph coloring (Stechly et al., 2023). Several studies claim that a bottleneck is in the feedback generation, and it is difficult to generate reliable feedback on their responses only by prompting themselves (Gou et al., 2024; Huang et al., 2024a; Olausson et al., 2024).
@@
Although our analysis of prior studies shows that intrinsic self-correction is difficult in general, some tasks have properties that make feedback generation easy and enable intrinsic self-correction. For example, CoVe (Dhuliawala et al., 2023) is an intrinsic self-correction method for tasks of generating multiple answers, such as Name some politicians who were born in NY, New York. Generated responses include multiple answers, but the feedback generation can be decomposed into easier sub-tasks of verifying each answer. Tasks with decomposable responses are one of the few groups of tasks for which verification is clearly easier than generation, which enables intrinsic self-correction. However, many real-world tasks do not satisfy this property.
@@
Self-Debug (2024d) Text-to-Code Code Interpreter
@@ 5.1 Self-Correction with External Tools or Knowledge
Given the observation that feedback generation is a bottleneck of self-correction (§4), improving feedback using external tools or knowledge is a promising direction. External tools used for self-correction include code interpreters for code generation tasks (Chen et al., 2024d; Gou et al., 2024) and symbolic reasoners for logical reasoning tasks (Pan et al., 2023). A popular source of knowledge is search engines, which are often used with queries generated from initial responses to retrieve information for validating their correctness (Gao et al., 2023; Jiang et al., 2023b). These prior studies widely agree that self-correction can improve LLM responses when reliable external tools or knowledge suitable for improving feedback are available.
@@
Although using external tools or knowledge is known to be effective in self-correction, we raise caution that the way of using external tools or knowledge influences the research questions we can verify (§3.1). As shown in Table 5, some prior studies (Gao et al., 2023; Yu et al., 2023; Zhao et al., 2023) use external knowledge only for self-correction, while they can also directly use external knowledge to improve the initial response generation process. For example, RARR (Gao et al., 2023) uses external knowledge to detect mistakes in initial responses, while it does not use any external knowledge when generating initial responses. These methods are reasonable when only focusing on [RQ3] the performance of final responses, but it is not fair to use them for evaluating [RQ2] whether self-correction can improve from the best-possible initial responses. In contrast, using code interpreters for self-correction (Gou et al., 2024; Chen et al., 2024d) can be regarded as using best-possible initial responses because there is no easy way to improve the initial response generation directly.
@@
Self-Edit (2023b) Code Generation Target Task GPT-3 (Code Executor and Test Cases) PyCodeGPT 110M Reference Code
Table 6: Self-correction with supervised fine-tuning. Most methods require large training datasets. “–” represents no fine-tuning.
@@ 5.2 Self-Correction with Fine-tuning
Prior work shows that fine-tuning LLMs for generating feedback or refining responses improves the self-correction capability. A common approach fine-tunes feedback models to generate reference feedback given initial responses and fine-tunes refinement models to generate reference answers given the initial responses and reference feedback (Ye et al., 2023; Lee et al., 2024; Saunders et al., 2022). Frameworks: The first approach fine-tunes the same model to correct its own responses. In this approach, most methods fine-tune models for all stages: initial responses, feedback, and refinement (Saunders et al., 2022; Ye et al., 2023; Lee et al., 2024). Another approach corrects responses from larger models using smaller fine-tuned models. This cross-model correction approach often instructs the larger models to refine their own responses using feedback from the smaller fine-tuned models (Yang et al., 2022b; Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024), which can be viewed as using the small fine-tuned models as external tools. Training Strategies: A popular approach is supervised fine-tuning, which fine-tunes self-correction modules on human-annotated feedback (Saunders et al., 2022), feedback from stronger models (Ye et al., 2023), or synthetic negative responses (Paul et al., 2024). As other approaches, to avoid the cost of collecting human feedback, self-corrective learning (Welleck et al., 2023) selects model-generated feedback that successfully refines responses as training data, and RL4L (Akyurek et al., 2023) uses reinforcement-learning. External Tools: Some works fine-tune models to refine responses given feedback from external tools. Self-Edit (Zhang et al., 2023b) uses the results on test cases evaluated by code executors for code generation, and Baldur (First et al., 2023) uses proof assistants for improving proof generation.
@@ 6 Strong Baselines
Self-correction involves multiple LLM calls for generating feedback and refinement. Therefore, to claim that [RQ3] the performance of the final outputs from self-correction frameworks is better than other approaches, it should be compared with sufficiently strong baselines, possibly relying on additional LLM calls or computational cost. Many self-correction studies do not compare their methods with strong baselines, although some studies pointed out this issue and compare self-correction with self-consistency (Gou et al., 2024; Huang et al., 2024a) or pass@k in code generation (Zhang et al., 2023b; Olausson et al., 2024). We encourage future research to compare self-correction with strong baselines, including self-consistency and generate-and-rank, to further explore RQ3.
@@
(Wang et al., 2023) is an approach that generates multiple responses for the same input and takes the majority vote of the final answers in reasoning tasks. The idea of selecting good responses using the consistency between multiple responses from the same model has also been extended to other tasks such as text generation (Manakul et al., 2023; Elaraby et al., 2023; Chen et al., 2024c) and code generation (Shi et al., 2022).
@@
is an approach that generates multiple responses and selects the best response using verifiers. Post-hoc approach ranks responses using self-evaluation (Weng et al., 2023; Zhang et al., 2023d), confidence (Manakul et al., 2023), fine-tuned verifiers (Cobbe et al., 2021; Shen et al., 2021; Lightman et al., 2024), or verifiers with external tools (Shi et al., 2022; Chen et al., 2023a; Ni et al., 2023). Feedback-guided decoding generates multiple responses and selects the best response for each reasoning step using generation probability (Hao et al., 2023; Tyen et al., 2024), prompted self-evaluation (Jung et al., 2022; Creswell and Shanahan, 2022; Xie et al., 2023; Yao et al., 2023; Miao et al., 2024), or fine-tuned verifiers (Uesato et al., 2022; Tafjord et al., 2022; Yang et al., 2022a; Asai et al., 2024).

### v3

@@ 1 Introduction
Self-correction is a popular approach to improve responses from large language models (LLMs) by refining them using LLMs during inference (Bai et al., 2022; Madaan et al., 2023). Extensive studies on self-correction have been conducted in various tasks, including arithmetic reasoning, code generation, and question answering (Gao et al., 2023; Shinn et al., 2023). The simplest approach of self-correction prompts LLMs to provide feedback on their own responses and refine the responses using the feedback (Huang et al., 2024a), under the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022). As in Figure 1, self-correction has also been studied using additional information for improving feedback, including external tools such as code interpreters (Chen et al., 2024e; Gou et al., 2024), external knowledge retrieved via web search (Gao et al., 2023; Jiang et al., 2023b), or fine-tuning (Welleck et al., 2023; Ye et al., 2023). However, recent studies also report negative results indicating that LLMs cannot self-correct (Huang et al., 2024a; Gou et al., 2024; Li et al., 2024b; Chen et al., 2024f) or even self-detect (Chen and Shu, 2024; Tyen et al., 2024; Hong et al., 2024; Jiang et al., 2024; Kamoi et al., 2024) their own mistakes at least in certain conditions. These conflicting observations indicate that further analysis of self-correction is needed.
@@
Timeframe.
This survey was originally published in June 2024 and primarily includes research papers and studies published up to and including May 2024. While papers published after this date are not comprehensively analyzed, they are briefly discussed in Section 12.
@@
Self-Debug (2024e) Fair-Asym. GPT-3.5-T, GPT-4 Code Interpreter Text-to-Code
@@
Tools , color=Red!100, fill=Purple!15, thick, text=black, l=5pt [ Fair , color=Black!100, fill=White!15, thick, text=black [ Code Interpreter or Compiler , color=Red!100, fill=Purple!15, thick, text=black [ Self-Debug (2024e), CRITIC (2024),
@@
Tools , color=Red!100, fill=Orange!15, thick, text=black, l=5pt [ Fair , color=Black!100, fill=White!15, thick, text=black [ Code Interpreter or Compiler , color=Red!100,, fill=Orange!15, thick, text=black, [ Self-Edit (2023a, SFT, PyCodeGPT) , leaf, color=Red!50, fill=Orange!15, thick, text=black ] ] [ Proof Assistant , color=Red!100, fill=Orange!15, thick, text=black, [ Baldur (2023, SFT) , leaf, color=Red!50, fill=Orange!15, thick, text=black ] ] ] ] ] ] [ Cross-Model
@@
Tools , color=Red!100, fill=Green!15, thick, text=black, l=5pt [ Code Interpreter or Compiler, color=Red!100, fill=Green!15, thick, text=black, text width=10em[ CodeRL (2022, RL), Self-Edit (2023a, SFT, GPT-3) , leaf, color=Red!50, fill=Green!15, thick, text=black, text width=20em] ] ] ] [ Multi-Agent
@@ 2.1 Frameworks
Cross-model correction generates feedback or refines the responses using models different from the model that generates initial responses. Cross-model correction has been mostly studied in the settings of correcting mistakes of large proprietary LLMs using small fine-tuned models (Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024) or multi-agent debate of multiple models with similar capabilities (Liang et al., 2024; Li et al., 2023; Cohen et al., 2023; Du et al., 2023; Zhang et al., 2024a; Chen et al., 2024b; Chan et al., 2024a; Wang et al., 2024a).
@@ 2.2 Sources of Feedback
Intrinsic self-correction prompts LLMs to generate feedback on their own responses. Prompting strategies include simple zero-shot or few-shot prompts (Madaan et al., 2023; Kim et al., 2023), decomposing the responses (Dhuliawala et al., 2024), and evaluating confidence (Varshney et al., 2023; Jiang et al., 2023b; Wu et al., 2024).
@@
Self-correction often relies on external information, including external tools such as code executors (Jiang et al., 2023a; Gou et al., 2024; Chen et al., 2024e; Stengel-Eskin et al., 2024), symbolic reasoners (Pan et al., 2023), proof assistant (First et al., 2023), or task-specific metrics (Xu et al., 2023), external knowledge from search engines (Jiang et al., 2023b; Gao et al., 2023; Zhao et al., 2023), Wikipedia (Yu et al., 2023; Zhao et al., 2023), or other corpora (Peng et al., 2023; Zhao et al., 2023), oracle information such as ground-truth answers (Kim et al., 2023; Shinn et al., 2023), human feedback (Chen et al., 2024a), or stronger models (Zhang et al., 2024c).
@@
Models fine-tuned for self-correction are another source of feedback, which are trained via supervised fine-tuning (Welleck et al., 2023; Ye et al., 2023; First et al., 2023; Paul et al., 2024; Han et al., 2024; Havrilla et al., 2024) or reinforcement learning (Le et al., 2022; Akyurek et al., 2023).
@@ 2.3 Tasks
Self-correction has been studied in various tasks, including Reasoning: arithmetic reasoning (Madaan et al., 2023; Nathani et al., 2023; Gou et al., 2024), code generation (Jiang et al., 2023a; Charalambous et al., 2023; Gou et al., 2024; Chen et al., 2024e; Olausson et al., 2024), proof generation (First et al., 2023), logical reasoning (Pan et al., 2023), Knowledge: closed-book QA (Shinn et al., 2023; Gao et al., 2023; Jiang et al., 2023b; Gou et al., 2024), Context-based Generation: dialogue generation (Madaan et al., 2023; Peng et al., 2023), text summarization (Saunders et al., 2022), Open-ended Generation: conditional text generation (Ye et al., 2023; Schick et al., 2023), story generation (Yang et al., 2022b), detoxification (Schick et al., 2021; Bai et al., 2022; Gou et al., 2024; Phute et al., 2024), Others: machine translation (Chen et al., 2024c; Raunak et al., 2023; Ki and Carpuat, 2024), information retrieval (Gero et al., 2023), vision language tasks (Yin et al., 2023; Ge et al., 2023; Zhou et al., 2024; Lee et al., 2024; Huang et al., 2024b; Liu et al., 2024), and prompt optimization (Pryzant et al., 2023; Mehrabi et al., 2024; Yang et al., 2024).
@@ 3.2 Frameworks for Verifying RQs
Realistic frameworks can be categorized by whether they use the best-possible initial responses. Fair self-correction represents frameworks that refine the best-possible initial responses. (1) Intrinsic self-correction (Huang et al., 2024a) uses the same model and information for initial response generation and self-correction. Intrinsic self-correction can be used to assess [RQ1] whether LLMs can self-correct based solely on their inherent capabilities. (2) Fair-asymmetric self-correction uses additional information for self-correction, but also uses information to improve initial response generation as much as possible. For example, self-correction with code interpreters (Chen et al., 2024e; Gou et al., 2024) is not intrinsic but fair because we cannot easily use code interpreters to directly improve the initial response generation. Fair-asymmetric self-correction can be used to evaluate [RQ2] whether LLMs can self-correct the best-possible initial responses using external information. Unfair self-correction (or unfair-asymmetric self-correction) represents frameworks that are practical but do not use the best-possible initial responses. For example, methods that use search engines only for self-correction (Gao et al., 2023; Yu et al., 2023) are unfair because they can use search engines to directly improve the initial response generation. Unfair self-correction can evaluate [RQ3] whether the final responses from self-correction outperform other methods but cannot evaluate [RQ2] whether self-correction can improve the best-possible initial responses.
@@ 4 Self-Correction with Prompting
However, recent studies report that intrinsic self-correction does not improve or even degrade the performance in tasks such as arithmetic reasoning, closed-book QA (Huang et al., 2024a; Gou et al., 2024), code generation (Gou et al., 2024; Olausson et al., 2024), plan generation (Valmeekam et al., 2023), and graph coloring (Stechly et al., 2023). Several studies claim that a bottleneck is in the feedback generation, and it is difficult to generate reliable feedback on their responses only by prompting themselves (Gou et al., 2024; Huang et al., 2024a; Olausson et al., 2024; Chen et al., 2024f).
@@
Although our analysis of prior studies shows that intrinsic self-correction is difficult in general, some tasks have properties that make feedback generation easy and enable intrinsic self-correction. For example, CoVe (Dhuliawala et al., 2024) is an intrinsic self-correction method for tasks of generating multiple answers, such as Name some politicians who were born in NY, New York. Generated responses include multiple answers, but the feedback generation can be decomposed into easier sub-tasks of verifying each answer. Tasks with decomposable responses are one of the few groups of tasks for which verification is clearly easier than generation, which enables intrinsic self-correction. However, many real-world tasks do not satisfy this property.
@@
Self-Debug (2024e) Text-to-Code Code Interpreter
@@ 5.1 Self-Correction with External Tools or Knowledge
Given the observation that feedback generation is a bottleneck of self-correction (§4), improving feedback using external tools or knowledge is a promising direction. External tools used for self-correction include code interpreters for code generation tasks (Chen et al., 2024e; Gou et al., 2024) and symbolic reasoners for logical reasoning tasks (Pan et al., 2023). A popular source of knowledge is search engines, which are often used with queries generated from initial responses to retrieve information for validating their correctness (Gao et al., 2023; Jiang et al., 2023b). These prior studies widely agree that self-correction can improve LLM responses when reliable external tools or knowledge suitable for improving feedback are available.
@@
Although using external tools or knowledge is known to be effective in self-correction, we raise caution that the way of using external tools or knowledge influences the research questions we can verify (§3.1). As shown in Table 5, some prior studies (Gao et al., 2023; Yu et al., 2023; Zhao et al., 2023) use external knowledge only for self-correction, while they can also directly use external knowledge to improve the initial response generation process. For example, RARR (Gao et al., 2023) uses external knowledge to detect mistakes in initial responses, while it does not use any external knowledge when generating initial responses. These methods are reasonable when only focusing on [RQ3] the performance of final responses, but it is not fair to use them for evaluating [RQ2] whether self-correction can improve from the best-possible initial responses. In contrast, using code interpreters for self-correction (Gou et al., 2024; Chen et al., 2024e) can be regarded as using best-possible initial responses because there is no easy way to improve the initial response generation directly.
@@
Self-Edit (2023a) Code Generation Target Task GPT-3 (Code Executor and Test Cases) PyCodeGPT 110M Reference Code
Table 6: Self-correction with supervised fine-tuning. Most methods require large training datasets. “–” on the “SFT Target” columns represents no fine-tuning.
@@ 5.2 Self-Correction with Fine-tuning
Prior work shows that fine-tuning LLMs for generating feedback or refining responses improves the self-correction capability. A common approach fine-tunes feedback models to generate reference feedback given initial responses and fine-tunes refinement models to generate reference answers given the initial responses and reference feedback (Ye et al., 2023; Lee et al., 2024; Saunders et al., 2022). Frameworks: The first approach fine-tunes the same model to correct its own responses. In this approach, most methods fine-tune models for all stages: initial responses, feedback, and refinement (Saunders et al., 2022; Ye et al., 2023; Lee et al., 2024). Another approach corrects responses from larger models using smaller fine-tuned models. This cross-model correction approach often instructs the larger models to refine their own responses using feedback from the smaller fine-tuned models (Yang et al., 2022b; Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024), which can be viewed as using the small fine-tuned models as external tools. Training Strategies: A popular approach is supervised fine-tuning, which fine-tunes self-correction modules on human-annotated feedback (Saunders et al., 2022), feedback from stronger models (Ye et al., 2023), or synthetic negative responses (Paul et al., 2024). As other approaches, to avoid the cost of collecting human feedback, self-corrective learning (Welleck et al., 2023) selects model-generated feedback that successfully refines responses as training data, GLoRe (Havrilla et al., 2024) constructs a synthetic refinement dataset using model-generated feedback, and RL4L (Akyurek et al., 2023) uses reinforcement-learning. External Tools: Some works fine-tune models to refine responses given feedback from external tools. Self-Edit (Zhang et al., 2023a) uses the results on test cases evaluated by code executors for code generation, and Baldur (First et al., 2023) uses proof assistants for improving proof generation.
@@ 6 Strong Baselines
Self-correction involves multiple LLM calls for generating feedback and refinement. Therefore, to claim that [RQ3] the performance of the final outputs from self-correction frameworks is better than other approaches, it should be compared with sufficiently strong baselines, possibly relying on additional LLM calls or computational cost. Many self-correction studies do not compare their methods with strong baselines, although some studies pointed out this issue and compare self-correction with self-consistency (Gou et al., 2024; Huang et al., 2024a) or pass@k in code generation (Zhang et al., 2023a; Olausson et al., 2024). We encourage future research to compare self-correction with strong baselines, including self-consistency and generate-and-rank, to further explore RQ3.
@@
(Wang et al., 2023) is an approach that generates multiple responses for the same input and takes the majority vote of the final answers in reasoning tasks. The idea of selecting good responses using the consistency between multiple responses from the same model has also been extended to other tasks such as text generation (Manakul et al., 2023; Elaraby et al., 2023; Chen et al., 2024d) and code generation (Shi et al., 2022).
@@
is an approach that generates multiple responses and selects the best response using verifiers. Post-hoc approach ranks responses using self-evaluation (Weng et al., 2023; Zhang et al., 2023b), confidence (Manakul et al., 2023), fine-tuned verifiers (Cobbe et al., 2021; Shen et al., 2021; Lightman et al., 2024), or verifiers with external tools (Shi et al., 2022; Chen et al., 2023; Ni et al., 2023). Feedback-guided decoding generates multiple responses and selects the best response for each reasoning step using generation probability (Hao et al., 2023; Tyen et al., 2024), prompted self-evaluation (Jung et al., 2022; Creswell and Shanahan, 2022; Xie et al., 2023; Yao et al., 2023; Miao et al., 2024), or fine-tuned verifiers (Uesato et al., 2022; Tafjord et al., 2022; Yang et al., 2022a; Asai et al., 2024).